OpenStudy (anonymous):

2sin^2(x)+sin(x)=0 solve on interval 0

myininaya (myininaya):

both terms have sin(x) in common so we can factor! \[\sin(x)(2\sin(x)+1)=0\]

myininaya (myininaya):

set both factors =0

myininaya (myininaya):

\[\sin(x)=0 ; 2 \sin(x)+1=0\]